OpenStudy (anonymous):

Use the given information to find the values of the 6 trig functions at the angle theta. Give exact answers. theta=arctan(-5/12) I know how to do it, but the part I am not too sure about is which one (x or y) does the negative sign belong to. I remember there was a way to figure it out, but I forgot, so can someone please help me refresh my memory? Thanks :)

OpenStudy (anonymous):

theta=-22.62derees

OpenStudy (anonymous):

To get exact values you want to draw the triangle associated with the angle.

OpenStudy (anonymous):

unfortunately we are not able to upload images right now, so i cant draw and show how its done =/

OpenStudy (anonymous):

now use sin(theta)=sin(-22.62)=-5/13,cos(tetha)=12/13, do he same to other trig functions

OpenStudy (anonymous):

Yeah, I can draw a triangle, but how do I know if it is x=-12 and y=5 or x=12 and y=-5?

OpenStudy (anonymous):

also csctheta=1/sintheta=-13/5, sectheta=1/cos theta=13/12

OpenStudy (anonymous):

and cot theta=1/tan theta=-12/5

OpenStudy (anonymous):

joemath is correct and you will understand them better if we draw them

OpenStudy (anonymous):

its kinda a poorly worded problem, usually it will tell you what quadrant it wants the angle to be in. I would just assume what mark.o said, let the angle be in the 4th quadrant. so it would be x = 12, y = -5

OpenStudy (anonymous):

Lol I'm sorry but I still don't understand how you guys came to conclusion that it is in the 4th quadrant... is it just a random guess? Or...? Sorry ><

OpenStudy (anonymous):

Because the problem doesnt specify which quadrant to go to, we are making a guess. Not really much we can do to get around that.

OpenStudy (anonymous):

Ok thanks~